### v1

# Emergence of a Fermi-surface in the current-driven Hidden state of 1T-TaS2

Abstract
We investigate the nature of the metallic metastable state in 1T-TaS2. Using microARPES, we measure the spatially-dependent modifications of the electronic structure of the sample following a short current pulse. We observe that, in some regions of the sample, a Fermi surface emerges, while other regions remain gapped. A detailed study of the band structure in these different regions suggests that the metallic parts are in a state similar to the nearly commensurate charge density wave (NC-CDW) state, where the gaps are suppressed and a band crosses the Fermi level. Furthermore, we find that the metallic and insulating regions of the sample exhibit different dispersions normal to the planes. This observation is consistent with a scenario in which the current pulse breaks the star-of-David dimers present in the commensurate charge density wave (C-CDW) state.
I Introduction
A metastable state refers to a state of a system that is not in its lowest energy state, but persists for a relatively long time before eventually transitioning to a more stable state. Some metastable states can be reached thermally (e.g., by quenching the system from an elevated temperature) Di Salvo et al. (1973); Sung et al. (2022), while others are ’hidden’ and can only be accessed through non-thermal perturbations Budden et al. (2021); Zhang et al. (2016); Zong et al. (2018); De La Torre et al. (2021).
One such hidden phase is the metallic state induced at low temperatures in 1T-TaS2 by various stimuli, including a short light pulse Stojchevska et al. (2014), charge transfer Ma et al. (2016), or a short electric current pulse Vaskivskyi et al. (2016). Regardless of the perturbation method, the hidden state exhibits a similar critical temperature Venturini et al. (2022) and relaxation time scales Vaskivskyi et al. (2015). Following the short pulse, the resistance of the samples is reduced by up to several orders of magnitude. This hidden metallic state has been shown to persist for hours and is fully reversible either when the sample is heated above a certain temperature Vaskivskyi et al. (2015) or subjected to an erasing pulse Stojchevska et al. (2014).
Due to the fast and controllable nature of the transition to the hidden state Maklar et al. (2023), it has been suggested for various applications, including programmable light manipulation Vaskivskyi et al. (2024) and charge configuration memory (CCM) devices Mihailovic et al. (2021); Mraz et al. (2022); Venturini et al. (2022); Vaskivskyi et al. (2016).
The origin of the insulating ground state of 1T-TaS2 is still not fully understood. The electronic-structure is governed by a charge density wave (CDW) instabilityBayliss et al. (1984); Rossnagel (2011). The main motif of the CDW is a rearrangement of every 13 Ta atoms in a star-of-David cluster where 12 Ta atoms are displaced towards the middle 13th atom. Below 352K domains of star-of-David super-cells start to form. The domains are separated by regions with no CDW and form an hexagonal lattice which is nearly commensurate (NC) with the atomic lattice Wu and Lieber (1989); Thomson et al. (1994); Wu and Lieber (1990). Below 183K all the star-of-David domains merge to create a commensurate CDW (C-CDW) state. The C-CDW vectors are rotated by 13.9 degrees from the atomic lattice vectors and form a new sized unit cell. The orientation of the CDW vectors breaks mirror symmetry, and gives rise to two degenerate configurations (see Fig.4) with a distinct chirality. The sample’s chirality is established during the formation of the CDW domains in the NC-CDW phase and persists into the C-CDW state Yang et al. (2022). It has been found that in the C-CDW state, these chiral domains can extend to millimeter-scale sizes Liu et al. (2023).
The phase transition from NC-CDW to C-CDW is marked by an abrupt increase in resistivity, followed by strong insulating behavior at lower temperatures. In the C-CDW state, the valence band splits into seven narrow bands, six of which are fully occupied, while the topmost band is half-filled. The presence of unpaired electrons suggests that C-CDW alone cannot fully explain the observed insulating state Tosatti and Fazekas (1976); Law and Lee (2017).
It has been proposed that the narrow half-filled band in 1T-TaS2 undergoes a Mott transition to an insulating stateWu et al. (2022); Fazekas and Tosatti (1980). In recent years, it has become evident that the out-of-plane stacking of star-of-David clusters plays a significant role in determining the electronic properties of 1T-TaS2Butler et al. (2020); Lee et al. (2019). Evidence for out-of-plane dimerization is observed during the transition to the C-CDW stateStahl et al. (2020), resulting in an even number of electrons per unit cell and leading to insulating behaviorRitschel et al. (2018). Understanding the nature of the metallic hidden state may provide insights into the origin of the insulating ground state.
In this paper, we use spatially resolved ARPES to track the evolution of the electronic structure of 1T-TaS2 after a short current pulse. Our main results are summarized in Fig. 1. Following a short current pulse, the resistance of the sample at low temperatures decreases by a factor of about four (Fig. 1a). In this state, we observe a non-uniform electronic structure, with different parts of the sample exhibiting distinct behaviors (Fig. 1b). Most of the sample remains in the C-CDW state, characterized by a full gap and a clear band folding due to the CDW both in-plane and out-of-plane. However, other regions, which contribute to the enhanced conductance in the hidden state, display a much weaker CDW effect and have a clear Fermi surface (Fig. 1c). The ARPES spectrum in these parts has some similarities with the NC-CDW spectrum although STM measurements find a completely different domain morphology in the two states Gerasimenko et al. (2019).
II Results and discussion
Narrow bridges were cut from a single crystal of 1T-TaS2 and four contacts were made using Ag epoxy (for more details about the sample preparation see the Methods section). In Fig. 1(a), we present the resistance as a function of temperature for a typical bridge. An abrupt change in resistance is observed when cooling into the C-CDW state at approximately 180K (blue line) and when warming back up at around 220K.
At a temperature of 40 K, a DC current pulse was applied, drastically lowering the sample’s resistance. This low resistance is maintained as long as the temperature remains below 70 K, which is consistent with the transition temperature observed when a short light pulse was used to drive the transition Vaskivskyi et al. (2015).
As higher current pulses are applied, the sample’s resistance decreases further, eventually saturating at around 140-190 mA for our typical sample dimensions. Upon warming, the resistance returns to its initial value. If the sample is cooled before a complete transition back to the C-CDW phase, it becomes locked in an intermediate state (see the purple line in Fig. 1a).
In Fig. 2 we show maps of the spectral weight at the Fermi-level over the entire cleaved area in the C-CDW state (Fig. 2(a)) and in the hidden state (2(b)). The color map represents the momentum-integrated intensity of the microARPES spectra at the Fermi-level. The spatial resolution is set by the spot size which is about 5 microns in this case. Inset of Fig. 2(a) shows a SEM scan of the measured sample. The cleaved area is marked.
In the C-CDW state, we observe a small, spatially uniform spectral weight at the Fermi level, originating from the tail of the flat band located at a binding energy of approximately 100 meV. In the hidden state, however, regions of high intensity indicate the metallic parts of the sample where the spectral gap is closed. The transition between the C-CDW and hidden states is fully reversible; in fact, the data in Fig. 2(a) was measured after the data in Fig. 2(b) by warming the sample and cooling it down again.
The use of a current pulse to drive the transition results in an inherently inhomogeneous state. Unlike a laser pulse, the current flow profile between the contacts cannot be precisely controlled and is generally non-uniform. Consequently, it is reasonable to assume that metallic regions form where the current density is sufficiently high. This inhomogeneity is not a characteristic of the hidden state but rather specific to the sample.
We observed different patterns of phase separation in the samples we measured and were unable to find defects on the sample surface that pin the metallic parts in the hidden state, suggesting that the switching is governed by the bulk current flow profile. Recently, it has been suggested that the metallic areas are more likely to form at the sample boundary Devidas et al. (2024).
In our study, we found an area of approximately 500 m2 that turned metallic, which is much larger than the typically reported domain size of 100 nm2 in the hidden state Gerasimenko et al. (2019). Each pixel in Fig. 2 contains many such domains.
Next, we compare in detail the spectra in the C-CDW state and the metallic regions of the hidden state. In Fig. 3, we present typical microARPES data along the direction, measured using 80 eV photon energy and a sub-micron spot size. Fig. 3(a) shows the timeline of the experiment. First, the sample was cooled to 40 K. The data before any current pulse was applied is shown in Fig. 3(b), revealing the signature of C-CDW gaps and a flat band around the point. A 140 mA, 1 ms long pulse was then applied, reducing the resistance by a factor of four. Fig. 3(c) presents the ARPES data after the current pulse, showing a suppression of the CDW gaps and a band crossing the Fermi level at k. The intensity at the point is significantly suppressed. The sample was then removed from the cryo-manipulator and allowed to warm up in the vacuum chamber for roughly 10 minutes before being cooled back to base temperature. Following this thermal cycle, the resistance partially recovered to about half of the original value. It is important to note that the exact temperature during this warm-up sequence is unknown. The data in Fig. 3(d) shows developed C-CDW gaps and a partial recovery of the flat band. A second current pulse was then applied and the resistance dropped to a tenth of the original resistance. Fig. 3(e) shows ARPES data after the second current pulse. Again, the C-CDW gaps are closed and a metallic band is formed. For comparison, we show in white dashed-line the band-dispersion measured at room temperature, in the NC-CDW state, using a 100 microns spot size. The agreement with the hidden state dispersion is remarkable.
In Fig. 3(f-h) we show energy distribution curves (EDCs) at k (f) , k (g) and at k (h) showing the flat band and the C-CDW gaps. At the point (f) we can see a suppression of the flat band in the hidden state. This band is not fully recovered after the thermal cycle. The small peak at is a signature of the folded band due to the C-CDW and is absent in the hidden phase. In (g) we find well developed gaps in the EDCs that correspond to the C-CDW state (marked with gray dashed lines), and a suppression of the gap in the EDCs measured in the hidden state. In (h) we show a metallic Fermi-edge in the EDCs taken in the hidden state and a gap at the Fermi-level for EDCs in the C-CDW state.
ARPES intensity maps measured in the hidden state at two different locations on the sample, one in a metallic region and the other in a gapped region are shown in Fig. 4. In the metallic part, at the Fermi-level, we find a clear Fermi-surface (Fig. 4a). This FS is the origin of the enhanced conductivity in the hidden state. The FS is similar to the FS found in the high temperature NC-CDW state Pillo et al. (1999). It consists of six elliptical pockets centered around the M-points. The Fermi pockets are gapped near the BZ edge because of the CDW, resulting in an incomplete Fermi surface Sohrt et al. (2014); Rossnagel (2011).
In the gapped regions, there is no Fermi surface, but we observe intensity at low binding energies extending almost to the Fermi level. Fig. 4(b) shows an intensity map at a binding energy of 30 meV. We find some spectral weight at the point of the atomic Brillouin zone (solid black lines) and weaker replicas at the points of the CDW-folded Brillouin zones (green dashed lines).
The electronic structure at higher binding energies indicates that a CDW deformation persists in the hidden state, even though the gaps are suppressed. In Fig. 4(c) and (d), we compare intensity maps at a binding energy of 300 meV in a metallic region (c) and a gapped region (d). Both datasets exhibit a chiral pattern attributed to mirror symmetry breaking caused by the formation of star-of-David clusters Yang et al. (2022).
Throughout the text, we refer to the band structure of 1T-TaS2 as chiral, in accordance with the accepted notation in the field. However, it is important to note that 1T-TaS2 is centrosymmetric, and therefore not chiral in the crystallographic sense. Instead, it is more accurately described as pseudochiral, a state characterized by the absence of mirror symmetry in the bulk 3D band structure Louat et al. (2024).
The two possible ways to arrange the BZ in the C-CDW are shown in Fig. 4(e). Interestingly, we find that the emerging Fermi surface also breaks mirror symmetry (see supplementary material), seemingly inheriting its chirality from the CDW domains.
Furthermore, by comparing the chiral patterns at the same metallic point in both the hidden state and before the application of the current pulse, we demonstrate that the chirality remains unchanged by the current pulse. This data is presented in the supplementary material. This strongly suggests that transitioning into the hidden state does not involve a complete melting of the CDW. This finding aligns with the distinct chirality of the domains observed in real-space STM data Gerasimenko et al. (2019).
The data from the metallic parts of the hidden state, including both the dispersion and the Fermi surface, closely resemble those of the NC-CDW statLahoud et al. (2014). However, STM measurements reveal that in the hidden phase, the C-CDW is split into a disordered network of domains, each approximately tens of nanometers in sizeGerasimenko et al. (2019). This domain configuration is distinct from that observed in the NC-CDW state.
The spot size used covers approximately 1000 domains. Given that the domain walls are around in widthWu and Lieber (1989), their contribution to the ARPES spectra is negligible. This, along with the distinct chirality of the Fermi surface, confirms that the metallic nature of the hidden state is not due to conducting domain walls.
After confirming that the CDW persists following the pulse current in both the metallic and gapped regions of the sample, we investigate the reasons for the differences between these regions. In Fig. 5, we present photon energy-dependent data measured at representative points in both regions. The out-of-plane dispersion was measured at a fixed parallel momentum, k, near the center of the reconstructed Brillouin zone, (as indicated by the dashed line in Fig. 5(a).
Fig. 5(a and b) correspond to a gapped region. Fig. 5(a) shows data along -M measured using 80eV photons and in (b) we show a color plot combining the EDCs at different kz values. The kz dispersion of the two uppermost sub-bands formed by the C-CDW, extracted by tracking the maxima in the EDCs, are shown in Fig. 5(e and f). The bandwidths of the two bands are 100 meV and 50 meV, respectively. Notably, both bands show a clear doubling of the expected periodicity, indicating a doubling of the unit cell. This observation is consistent with a Q=0.5 charge density wave (CDW) out-of-plane or a dimerization of pairs of star-of-David clusters on two neighboring planes.
When examining the kz dispersion in the metallic regions, we observe a different behavior. In Figs. 5(c) and 5(d), we present data along -M measured using 80eV photons and the photon energy dependence of the EDCs measured in a metallic part of the sample. The extracted dispersion is shown in Fig. 5(g), revealing only one band at this k∥ value. The bandwidth is approximately 200 meV, which is significantly larger than that observed in the insulating region. Notably, we do not find double periodicity; instead, the dispersion is in agreement with the atomic lattice unit cell.
This can explain the metallic behavior. In the C-CDW state, the doubling of the unit cell perpendicular to the planes results in two electrons per unit cell, causing the valence band to be fully occupied, as expected in a band insulator. The current pulse disrupts the dimerization out-of-plane but maintains the CDW deformation within the planes. Although the in-plane hopping compared to the on-site Coulomb interaction places the sample beyond the Mott limit, the significant out-of-plane dispersion is sufficient to restore the metallic behavior of the 3D sample.
III Summary
Our study reveals that the hidden state in 1T-TaS2 exhibits a hole-like band crossing the Fermi level, resulting in a Fermi surface reminiscent of the NC-CDW phase. This similarity is not unexpected, given the structural resemblance between the two phases, both characterized by domains of commensurate CDW. However, the underlying mechanisms rendering these phases metallic remain elusive. While it was hypothesized that conducting domain walls, where the order parameter , contribute to the conductivity by forming channels, our ARPES measurements of the hidden phase suggest a different scenario.
We demonstrate that the hidden Fermi surface displays distinct chirality, originating from the C-CDW arrangements—a property inherent to the domains rather than the domain walls. Additionally, ARPES, which is sensitive to the surface area, indicates that the contribution from domain walls is negligible. This observation implies that each domain is intrinsically metallic and possesses an in-plane C-CDW configuration, suggesting that additional factors influence the metallic behavior.
We propose that the metallic behavior originates from the out-of-plane electronic dispersion. The periodicity of the dispersion along k⟂ indicates a doubling of the Brillouin zone in the out-of-plane direction within the C-CDW state, consistent with the previously reported layer dimerization Stahl et al. (2020). Upon transitioning to the hidden state, the out-of-plane dispersion changes, suggesting that star-of-David clusters on neighboring planes are no longer correlated. This leads to a profound change in the electronic structure: instead of having two electrons per unit cell in the C-CDW state, we now have a metallic half-filled band. The resulting band structure is unusual for a TMD, as it has a larger hopping constant normal to the Ta planes than within the planes.
IV Methods
IV.1 Sample preparation
1T-TaS2 single crystal were prepared as described in Ref. Ribak et al. (2017). Samples of 1T-TaS2 were prepared for nanoARPES experiments. Single crystals were cut into narrow bridges 200-500 in width. The samples were glued to a sapphire substrate and 4 contacts were made using Ag epoxy. The samples were cooled to 40K and cleaved under UHV conditions. The 4 contacts were used to drive a DC current pulse and to monitor the sample resistance. We emphasis that during the ARPES measurement no current was applied to the sample.
IV.2 ARPES
Spatially resolved ARPES experiments were carried out in the I05 beamline at Diamond light source, UK. The spatial resolution is set by the beam spot-size, which in our experiment was either 500nm using a Zone-plate or 5 micrometers using a capillary mirror. Energy resolution was set to about 50meV. Data in the NC-CDW state was measured at the CASSIOPEE beamline at Soleil, France. Measurements were done at room temperature using 80eV photon energy. Beam spot size was 100 microns. Energy resolution was set to 20meV. The Fermi-level was determined by measuring a metallic sample at the same conditions.
IV.3 calculation of k⟂
We have used the three-step-model for the out-of-plane momentum k Damascelli (2004), where is the inner potential. We have used , obtained by following the normal-crystal periodicity over 3 BZs, demanding ().
V Acknowledgments
We thank Assa Auerbach and Daniel Podolsky for useful discussions. We acknowledge Diamond Light Source for time on Beamline i05 under Proposal SI33131. We acknowledge SOLEIL for provision of synchrotron radiation facilities under proposal 20230771 and we would like to thank Chiara Bigi and Francois Bertran for assistance in using beamline CASSIOPEE. The work at the Technion was supported by Israeli Science Foundation grant number ISF-1263/21.

### v2

@@
 Abstract
-We investigate the nature of the metallic metastable state in 1T-TaS2.
-Using microARPES, we measure the spatially-dependent modifications of the electronic structure of the sample following a short current pulse.
-We observe that, in some regions of the sample, a Fermi surface emerges, while other regions remain gapped.
-A detailed study of the band structure in these different regions suggests that the metallic parts are in a state similar to the nearly commensurate charge density wave (NC-CDW) state, where the gaps are suppressed and a band crosses the Fermi level.
-Furthermore, we find that the metallic and insulating regions of the sample exhibit different dispersions normal to the planes.
-This observation is consistent with a scenario in which the current pulse breaks the star-of-David dimers present in the commensurate charge density wave (C-CDW) state.
-I Introduction
-A metastable state refers to a state of a system that is not in its lowest energy state, but persists for a relatively long time before eventually transitioning to a more stable state.
-Some metastable states can be reached thermally (e.g., by quenching the system from an elevated temperature) Di Salvo et al.
-(1973); Sung et al.
-(2022), while others are ’hidden’ and can only be accessed through non-thermal perturbations Budden et al.
-(2021); Zhang et al.
-(2016); Zong et al.
-(2018); De La Torre et al.
-(2021).
-One such hidden phase is the metallic state induced at low temperatures in 1T-TaS2 by various stimuli, including a short light pulse Stojchevska et al.
-(2014), charge transfer Ma et al.
-(2016), or a short electric current pulse Vaskivskyi et al.
-(2016).
-Regardless of the perturbation method, the hidden state exhibits a similar critical temperature Venturini et al.
-(2022) and relaxation time scales Vaskivskyi et al.
-(2015).
+The origin of the insulating state in 1T-TaS2 has long been a subject of debate.
+A short current pulse transforms this insulating state into a metastable metallic phase.
+Using micro-ARPES, we investigate the electronic structure of this phase and uncover spatially dependent modifications caused by the current pulse.
+In some regions of the sample, a Fermi surface emerges, while others remain gapped.
+Detailed band structure analysis reveals that the metallic regions exhibit an electronic structure similar to that observed in the high-temperature phase of 1T-TaS2, characterized by suppressed energy gaps and bands crossing the Fermi level.
+Furthermore, the metallic and insulating regions display distinct dispersions along the out-of-plane direction.
+These observations suggest a scenario in which the current pulse breaks the star-of-David dimers characteristic of the insulating phase, implicating these dimers as the likely origin of the insulating behavior in 1T-TaS2.
+Metastable and hidden states have emerged as key areas of interest in condensed matter physics, offering insights into novel phases of matter and exotic physical properties.
+They also present promising avenues for technological applications.
+A metastable state occurs when a system becomes trapped in a local energy minimum rather than reaching its global ground state.
+These states typically arise from competing orders[1, 2, 3], topological defects[4], structural configurations[5], and other degrees of freedom that create a complex energy landscape with multiple stable configurations.
+Some metastable states can be accessed thermally, for example, by rapidly cooling a material from elevated temperatures [6, 7].
+Distinctly, hidden states represent a subset of metastable states that require non-thermal stimuli.
+These hidden configurations can only be reached through external drivers like ultrafast laser pulses or strong electric fields.
+[8, 9, 10, 11].
+The study of these metastable states provides valuable insights into material properties and underlying physical mechanisms.
+By examining metastable configurations, we can better understand fundamental interactions in materials, for example electron-phonon coupling[12] and magnetic-lattice interactions[9].
+One such hidden phase is the metallic state induced at low temperatures in 1T-TaS2 by various stimuli, including a short light pulse [13], charge transfer [14], or a short electric current pulse [15].
+Regardless of the perturbation method, the hidden state exhibits a similar critical temperature [16] and relaxation time scales [17].
 Following the short pulse, the resistance of the samples is reduced by up to several orders of magnitude.
-This hidden metallic state has been shown to persist for hours and is fully reversible either when the sample is heated above a certain temperature Vaskivskyi et al.
-(2015) or subjected to an erasing pulse Stojchevska et al.
-(2014).
-Due to the fast and controllable nature of the transition to the hidden state Maklar et al.
-(2023), it has been suggested for various applications, including programmable light manipulation Vaskivskyi et al.
-(2024) and charge configuration memory (CCM) devices Mihailovic et al.
-(2021); Mraz et al.
-(2022); Venturini et al.
-(2022); Vaskivskyi et al.
-(2016).
-The origin of the insulating ground state of 1T-TaS2 is still not fully understood.
-The electronic-structure is governed by a charge density wave (CDW) instabilityBayliss et al.
-(1984); Rossnagel (2011).
+This hidden metallic state has been shown to persist for hours and is fully reversible either when the sample is heated above a certain temperature [17] or subjected to an erasing pulse [13].
+Due to the fast and controllable nature of the transition to the hidden state [18], it has been suggested for various applications, including programmable light manipulation [19] and charge configuration memory (CCM) devices [20, 21, 16, 15].
+1T-TaS2 exhibits an insulating ground state, whose origin remains unknown.
+The electronic-structure is governed by a charge density wave (CDW) instability [22, 23].
 The main motif of the CDW is a rearrangement of every 13 Ta atoms in a star-of-David cluster where 12 Ta atoms are displaced towards the middle 13th atom.
-Below 352K domains of star-of-David super-cells start to form.
-The domains are separated by regions with no CDW and form an hexagonal lattice which is nearly commensurate (NC) with the atomic lattice Wu and Lieber (1989); Thomson et al.
-(1994); Wu and Lieber (1990).
-Below 183K all the star-of-David domains merge to create a commensurate CDW (C-CDW) state.
+Below domains of star-of-David super-cells start to form.
+The domains are separated by regions with no CDW and form a hexagonal lattice which is nearly commensurate (NC) with the atomic lattice [24, 25, 26].
+Below all the star-of-David domains merge to create a commensurate CDW (C-CDW) state.
 The C-CDW vectors are rotated by 13.9 degrees from the atomic lattice vectors and form a new sized unit cell.
-The orientation of the CDW vectors breaks mirror symmetry, and gives rise to two degenerate configurations (see Fig.4) with a distinct chirality.
-The sample’s chirality is established during the formation of the CDW domains in the NC-CDW phase and persists into the C-CDW state Yang et al.
-(2022).
-It has been found that in the C-CDW state, these chiral domains can extend to millimeter-scale sizes Liu et al.
-(2023).
+The orientation of the CDW vectors breaks mirror symmetry, and gives rise to two degenerate configurations (see Fig.
+4E) with a distinct chirality.
+The sample’s chirality is established during the formation of the CDW domains in the NC-CDW phase and persists into the C-CDW state [27], where these chiral domains can extend to millimeter-scale sizes [28].
 The phase transition from NC-CDW to C-CDW is marked by an abrupt increase in resistivity, followed by strong insulating behavior at lower temperatures.
 In the C-CDW state, the valence band splits into seven narrow bands, six of which are fully occupied, while the topmost band is half-filled.
-The presence of unpaired electrons suggests that C-CDW alone cannot fully explain the observed insulating state Tosatti and Fazekas (1976); Law and Lee (2017).
-It has been proposed that the narrow half-filled band in 1T-TaS2 undergoes a Mott transition to an insulating stateWu et al.
-(2022); Fazekas and Tosatti (1980).
-In recent years, it has become evident that the out-of-plane stacking of star-of-David clusters plays a significant role in determining the electronic properties of 1T-TaS2Butler et al.
-(2020); Lee et al.
-(2019).
-Evidence for out-of-plane dimerization is observed during the transition to the C-CDW stateStahl et al.
-(2020), resulting in an even number of electrons per unit cell and leading to insulating behaviorRitschel et al.
-(2018).
+The presence of unpaired electrons suggests that the C-CDW alone cannot fully explain the observed insulating state [29, 30].
+It has been proposed that the narrow half-filled band in 1T-TaS2 undergoes a Mott transition to an insulating state [31, 32].
+In recent years, it has become evident that the out-of-plane stacking of star-of-David clusters plays a significant role in determining the electronic properties of 1T-TaS2 [33, 34].
+Evidence for out-of-plane dimerization is observed during the transition to the C-CDW state [35], resulting in an even number of electrons per unit cell and leading to insulating behavior [36].
+STM measurement reveal that the monolayer 1T-TaS2 is insulating[37], consistent with Mott insulating behavior in the absence of inter-layer coupling.
+However, it remains unclear whether this behavior persists in the bulk, where out-of-plane interactions play a critical role.
 Understanding the nature of the metallic hidden state may provide insights into the origin of the insulating ground state.
@@
 1B).
-Most of the sample remains in the C-CDW state, characterized by a full gap and a clear band folding due to the CDW both in-plane and out-of-plane.
+Most of the sample remains in the C-CDW state, characterized by a full gap and a clear band folding due to the CDW, both in-plane and out-of-plane.
 However, other regions, which contribute to the enhanced conductance in the hidden state, display a much weaker CDW effect and have a clear Fermi surface (Fig.
 1C).
-The ARPES spectrum in these parts has some similarities with the NC-CDW spectrum although STM measurements find a completely different domain morphology in the two states Gerasimenko et al.
-(2019).
-II Results and discussion
+The ARPES spectrum in these parts has some similarities with the NC-CDW spectrum although STM measurements find a completely different domain morphology in the two states [4].
+Results
+Experimental setup
 Narrow bridges were cut from a single crystal of 1T-TaS2 and four contacts were made using Ag epoxy (for more details about the sample preparation see the Methods section).
 In Fig.
-1(a), we present the resistance as a function of temperature for a typical bridge.
-An abrupt change in resistance is observed when cooling into the C-CDW state at approximately 180K (blue line) and when warming back up at around 220K.
-At a temperature of 40 K, a DC current pulse was applied, drastically lowering the sample’s resistance.
-This low resistance is maintained as long as the temperature remains below 70 K, which is consistent with the transition temperature observed when a short light pulse was used to drive the transition Vaskivskyi et al.
-(2015).
+1A, we present the resistance as a function of temperature for a typical bridge.
+An abrupt change in resistance is observed when cooling into the C-CDW state at approximately (blue line) and when warming back up at approximately (red line).
+At a temperature of , a current pulse was applied, drastically lowering the sample’s resistance.
+This low resistance is maintained as long as the temperature remains below 70 K, which is consistent with the transition temperature observed when a short light pulse was used to drive the transition [17].
 As higher current pulses are applied, the sample’s resistance decreases further, eventually saturating at around 140-190 mA for our typical sample dimensions.
@@
 1A).
+Spatial distribution of the hidden state
 In Fig.
 2 we show maps of the spectral weight at the Fermi-level over the entire cleaved area in the C-CDW state (Fig.
-2(a)) and in the hidden state (2(b)).
+2A) and in the hidden state (Fig.
+2B).
 The color map represents the momentum-integrated intensity of the microARPES spectra at the Fermi-level.
 The spatial resolution is set by the spot size which is about 5 microns in this case.
-Inset of Fig.
-2(a) shows a SEM scan of the measured sample.
+The inset of Fig.
+2A shows an SEM image of the measured sample.
 The cleaved area is marked.
@@
 The transition between the C-CDW and hidden states is fully reversible; in fact, the data in Fig.
-2(a) was measured after the data in Fig.
-2(b) by warming the sample and cooling it down again.
+2A was measured after the data in Fig.
+2B by warming the sample and cooling it down again.
 The use of a current pulse to drive the transition results in an inherently inhomogeneous state.
@@
 We observed different patterns of phase separation in the samples we measured and were unable to find defects on the sample surface that pin the metallic parts in the hidden state, suggesting that the switching is governed by the bulk current flow profile.
-Recently, it has been suggested that the metallic areas are more likely to form at the sample boundary Devidas et al.
-(2024).
-In our study, we found an area of approximately 500 m2 that turned metallic, which is much larger than the typically reported domain size of 100 nm2 in the hidden state Gerasimenko et al.
-(2019).
-Each pixel in Fig.
-2 contains many such domains.
+Recently, it has been suggested that the metallic areas are more likely to form at the sample boundary [38].
+Band dispersion of the hidden state
 Next, we compare in detail the spectra in the C-CDW state and the metallic regions of the hidden state.
 In Fig.
-3, we present typical microARPES data along the direction, measured using 80 eV photon energy and a sub-micron spot size.
+3, we present typical microARPES data along the direction, measured using photon energy and a sub-micron spot size.
 Fig.
-3(a) shows the timeline of the experiment.
-First, the sample was cooled to 40 K.
+3A shows the timeline of the experiment.
+First, the sample was cooled to .
 The data before any current pulse was applied is shown in Fig.
-3(b), revealing the signature of C-CDW gaps and a flat band around the point.
+3B, revealing the signature C-CDW gaps and a flat band around the point.
 A 140 mA, 1 ms long pulse was then applied, reducing the resistance by a factor of four.
 Fig.
-3(c) presents the ARPES data after the current pulse, showing a suppression of the CDW gaps and a band crossing the Fermi level at k.
-The intensity at the point is significantly suppressed.
-The sample was then removed from the cryo-manipulator and allowed to warm up in the vacuum chamber for roughly 10 minutes before being cooled back to base temperature.
+3C presents the ARPES data after the current pulse, showing a suppression of the CDW gaps and a band crossing the Fermi level at k.
+The intensity of the flat band at the point is significantly suppressed.
+The sample was then removed from the cryo-manipulator and allowed to warm up in the vacuum chamber for approximately 10 minutes before being cooled back to base temperature.
 Following this thermal cycle, the resistance partially recovered to about half of the original value.
@@
 The data in Fig.
-3(d) shows developed C-CDW gaps and a partial recovery of the flat band.
+3D show developed C-CDW gaps and a partial recovery of the flat band.
 A second current pulse was then applied and the resistance dropped to a tenth of the original resistance.
 Fig.
-3(e) shows ARPES data after the second current pulse.
+3E shows ARPES data after the second current pulse.
 Again, the C-CDW gaps are closed and a metallic band is formed.
@@
 In Fig.
-3(f-h) we show energy distribution curves (EDCs) at k (f) , k (g) and at k (h) showing the flat band and the C-CDW gaps.
+3F-H we show energy distribution curves (EDCs) at k (F) , k (G) and at k (H), showing the flat band and the C-CDW gaps.
 At the point (F) we can see a suppression of the flat band in the hidden state.
-This band is not fully recovered after the thermal cycle.
+The band does not fully recover after the thermal cycle.
 The small peak at is a signature of the folded band due to the C-CDW and is absent in the hidden phase.
@@
 In (H) we show a metallic Fermi-edge in the EDCs taken in the hidden state and a gap at the Fermi-level for EDCs in the C-CDW state.
+Chirality of the Fermi surface
 ARPES intensity maps measured in the hidden state at two different locations on the sample, one in a metallic region and the other in a gapped region are shown in Fig.
@@
 This FS is the origin of the enhanced conductivity in the hidden state.
-The FS is similar to the FS found in the high temperature NC-CDW state Pillo et al.
-(1999).
+The FS is similar to the FS found in the high temperature NC-CDW state [39].
 It consists of six elliptical pockets centered around the M-points.
-The Fermi pockets are gapped near the BZ edge because of the CDW, resulting in an incomplete Fermi surface Sohrt et al.
-(2014); Rossnagel (2011).
+The Fermi pockets are gapped near the BZ edges, resulting in an incomplete Fermi surface [40, 23].
 In the gapped regions, there is no Fermi surface, but we observe intensity at low binding energies extending almost to the Fermi level.
 Fig.
-4(b) shows an intensity map at a binding energy of 30 meV.
+4B shows an intensity map at a binding energy of .
 We find some spectral weight at the point of the atomic Brillouin zone (solid black lines) and weaker replicas at the points of the CDW-folded Brillouin zones (green dashed lines).
@@
 In Fig.
-4(c) and (d), we compare intensity maps at a binding energy of 300 meV in a metallic region (c) and a gapped region (d).
-Both datasets exhibit a chiral pattern attributed to mirror symmetry breaking caused by the formation of star-of-David clusters Yang et al.
-(2022).
+4C and D, we compare intensity maps at a binding energy of 300 meV in a metallic region (C) and a gapped region (D).
+Both datasets exhibit a chiral pattern attributed to mirror symmetry breaking caused by the formation of star-of-David clusters [27].
+We emphasize that no measurable differences were observed between the gapped regions of the hidden state and the C-CDW spectra (see figure S1).
 Throughout the text, we refer to the band structure of 1T-TaS2 as chiral, in accordance with the accepted notation in the field.
 However, it is important to note that 1T-TaS2 is centrosymmetric, and therefore not chiral in the crystallographic sense.
-Instead, it is more accurately described as pseudochiral, a state characterized by the absence of mirror symmetry in the bulk 3D band structure Louat et al.
-(2024).
-The two possible ways to arrange the BZ in the C-CDW are shown in Fig.
-4(e).
-Interestingly, we find that the emerging Fermi surface also breaks mirror symmetry (see supplementary material), seemingly inheriting its chirality from the CDW domains.
-Furthermore, by comparing the chiral patterns at the same metallic point in both the hidden state and before the application of the current pulse, we demonstrate that the chirality remains unchanged by the current pulse.
-This data is presented in the supplementary material.
-This strongly suggests that transitioning into the hidden state does not involve a complete melting of the CDW.
-This finding aligns with the distinct chirality of the domains observed in real-space STM data Gerasimenko et al.
-(2019).
-The data from the metallic parts of the hidden state, including both the dispersion and the Fermi surface, closely resemble those of the NC-CDW statLahoud et al.
-(2014).
-However, STM measurements reveal that in the hidden phase, the C-CDW is split into a disordered network of domains, each approximately tens of nanometers in sizeGerasimenko et al.
-(2019).
+Instead, it is more accurately described as pseudochiral, a state characterized by the absence of mirror symmetry in the bulk 3D band structure [41].
+The two possible configurations of the BZ in the C-CDW phase are illustrated in Fig.
+4E.
+Remarkably, our results indicate that the resulting Fermi surface also breaks mirror symmetry, seemingly inheriting its chirality from the underlying CDW domains.
+Moreover, by comparing the chiral patterns at the same metallic point both in the hidden state and prior to the application of the current pulse, we show that the chirality remains unchanged by the current pulse (see figure S2).
+This evidence, shown in the supplementary material, strongly suggests that the transition into the hidden state does not involve a complete melting of the CDW.
+This conclusion is in agreement with the distinct chirality of domains observed in real-space STM data [4].
+The data from the metallic parts of the hidden state, both the dispersion and the Fermi surface, closely resemble those of the NC-CDW stat [42].
+However, STM measurements reveal that in the hidden phase, the C-CDW is fragmented into a disordered network of domains, each approximately tens of nanometers in size [4].
 This domain configuration is distinct from that observed in the NC-CDW state.
-The spot size used covers approximately 1000 domains.
-Given that the domain walls are around in widthWu and Lieber (1989), their contribution to the ARPES spectra is negligible.
-This, along with the distinct chirality of the Fermi surface, confirms that the metallic nature of the hidden state is not due to conducting domain walls.
-After confirming that the CDW persists following the pulse current in both the metallic and gapped regions of the sample, we investigate the reasons for the differences between these regions.
+The spot size used in our experiment covers approximately 1000 domains.
+Given that the domain walls are around in width [24], their contribution to the ARPES spectra is negligible.
+Combined with the distinct chirality of the Fermi surface we find(see figure S3), this confirms that the metallic nature of the hidden state is not due to conducting domain walls.
+Out-of-plane dispersion measurements
+After confirming that the CDW persists in both the metallic and gapped regions of the sample following the current pulse, we turn our attention to understanding the differences between these regions.
 In Fig.
-5, we present photon energy-dependent data measured at representative points in both regions.
-The out-of-plane dispersion was measured at a fixed parallel momentum, k, near the center of the reconstructed Brillouin zone, (as indicated by the dashed line in Fig.
-5(a).
+5, we present photon-energy dependent data measured at representative points in both areas.
+The out-of-plane dispersion was recorded at a fixed parallel momentum, k (indicated by the red dot in the inset of Fig.
+5D).
+This point lies near the center of the reconstructed Brillouin zone, , as seen by the faint replica of the band at a binding energy of .
+Panels 5(A and B) correspond to a gapped region while panels 5C and D are from a metallic region.
 Fig.
-5(a and b) correspond to a gapped region.
-Fig.
-5(a) shows data along -M measured using 80eV photons and in (b) we show a color plot combining the EDCs at different kz values.
-The kz dispersion of the two uppermost sub-bands formed by the C-CDW, extracted by tracking the maxima in the EDCs, are shown in Fig.
-5(e and f).
-The bandwidths of the two bands are 100 meV and 50 meV, respectively.
+5A and C show data along -M measured using 77 eV photons in the insulating and metallic regions, respectively.
+The white dashed lines represent the in-plane momentum at which the dispersion out-of-plane was measured.
+In Fig.
+5B and D we show a color plot combining the EDCs at different kz values.
+The kz dispersion of the two uppermost sub-bands formed by the C-CDW, extracted by tracking the maxima in the EDCs, are shown in Fig. 5E and F.
+The bandwidths of the two bands are approximately 100 meV and 50 meV, respectively.
 Notably, both bands show a clear doubling of the expected periodicity, indicating a doubling of the unit cell.
-This observation is consistent with a Q=0.5 charge density wave (CDW) out-of-plane or a dimerization of pairs of star-of-David clusters on two neighboring planes.
+This observation is consistent with a charge density wave out-of-plane with two unit-cell periodicity or a dimerization of pairs of star-of-David clusters on two neighboring planes.
+The CDW replicas exhibit significantly weaker intensity compared to the main band.
+The double periodicity becomes more pronounced when the dispersion is measured at the point instead of the point.
+The dashed white line in Fig.5B represents the results of a DFT calculation for an alternating stacking t02 configuration, where two vertically aligned star-of-David clusters form pairs within each bilayer, and these pairs are shifted relative to the pairs in neighboring bilayers [36].
+Our results show qualitative agreement with the DFT calculation.
+We observe a clear doubling of the unit cell along the out-of-plane direction, with the spectrum gapped at all kz values.
+Previous studies were somewhat inconclusive on this point [43, 44].
+The difference could be related to the significantly smaller light spot used in our study.
 When examining the kz dispersion in the metallic regions, we observe a different behavior.
 In Figs.
-5(c) and 5(d), we present data along -M measured using 80eV photons and the photon energy dependence of the EDCs measured in a metallic part of the sample.
+5C and 5D, we present data along -M measured using photons, and photon energy dependence of the EDCs measured in a metallic part of the sample, respectivley.
 The extracted dispersion is shown in Fig.
-5(g), revealing only one band at this k∥ value.
+5G, revealing only one band at this k∥ value.
 The bandwidth is approximately 200 meV, which is significantly larger than that observed in the insulating region.
-Notably, we do not find double periodicity; instead, the dispersion is in agreement with the atomic lattice unit cell.
-This can explain the metallic behavior.
-In the C-CDW state, the doubling of the unit cell perpendicular to the planes results in two electrons per unit cell, causing the valence band to be fully occupied, as expected in a band insulator.
-The current pulse disrupts the dimerization out-of-plane but maintains the CDW deformation within the planes.
-Although the in-plane hopping compared to the on-site Coulomb interaction places the sample beyond the Mott limit, the significant out-of-plane dispersion is sufficient to restore the metallic behavior of the 3D sample.
-III Summary
-Our study reveals that the hidden state in 1T-TaS2 exhibits a hole-like band crossing the Fermi level, resulting in a Fermi surface reminiscent of the NC-CDW phase.
-This similarity is not unexpected, given the structural resemblance between the two phases, both characterized by domains of commensurate CDW.
-However, the underlying mechanisms rendering these phases metallic remain elusive.
-While it was hypothesized that conducting domain walls, where the order parameter , contribute to the conductivity by forming channels, our ARPES measurements of the hidden phase suggest a different scenario.
-We demonstrate that the hidden Fermi surface displays distinct chirality, originating from the C-CDW arrangements—a property inherent to the domains rather than the domain walls.
-Additionally, ARPES, which is sensitive to the surface area, indicates that the contribution from domain walls is negligible.
-This observation implies that each domain is intrinsically metallic and possesses an in-plane C-CDW configuration, suggesting that additional factors influence the metallic behavior.
-We propose that the metallic behavior originates from the out-of-plane electronic dispersion.
-The periodicity of the dispersion along k⟂ indicates a doubling of the Brillouin zone in the out-of-plane direction within the C-CDW state, consistent with the previously reported layer dimerization Stahl et al.
-(2020).
-Upon transitioning to the hidden state, the out-of-plane dispersion changes, suggesting that star-of-David clusters on neighboring planes are no longer correlated.
-This leads to a profound change in the electronic structure: instead of having two electrons per unit cell in the C-CDW state, we now have a metallic half-filled band.
-The resulting band structure is unusual for a TMD, as it has a larger hopping constant normal to the Ta planes than within the planes.
-IV Methods
-IV.1 Sample preparation
-1T-TaS2 single crystal were prepared as described in Ref.
-Ribak et al.
-(2017).
-Samples of 1T-TaS2 were prepared for nanoARPES experiments.
-Single crystals were cut into narrow bridges 200-500 in width.
+Notably, we do not find the double periodicity; instead, the dispersion is in agreement with the atomic lattice unit cell.
+Discussion
+Our study reveals that the hidden state in 1T-TaS2 hosts a hole-like band crossing the Fermi level, resulting in a Fermi surface reminiscent of the nearly commensurate NC-CDW phase.
+This similarity is not surprising, given the structural resemblance between the two phases—both characterized by domains of commensurate CDW order.
+However, the mechanisms underlying the metallicity of these phases remain elusive.
+While it has been proposed that conducting domain walls—regions where the order parameter - form channels that contribute to the overall conductivity, our ARPES measurements suggest that this mechanism alone cannot account for the observed metallic behavior.
+Although such channels may be locally conductive, they are insufficient to explain the emergence of a robust Fermi surface and substantial spectral weight at the Fermi level.
+We show that the hidden Fermi surface exhibits distinct chirality arising from the C-CDW domain arrangements - an intrinsic property of the domains themselves rather than the domain walls.
+Furthermore, the significant spectral weight observed in ARPES, which is surface-sensitive, indicates that each domain is intrinsically metallic, and the contribution from domain walls is minimal.
+Importantly, the in-plane CDW motif remains essentially intact across both insulating and metallic regions, suggesting that the metallicity does not result from CDW melting or disordering, but from a more subtle structural transformation.
+Our findings establish a compelling link between the stacking configuration of 1T-TaS2 and its emergent electronic properties.
+While the monolayer is likely a Mott insulator [37], in the bulk, first-principles calculations have shown that the out-of-plane stacking of star-of-David clusters plays a decisive role in shaping the in-plane electronic dispersion, modulating gap structures and bandwidths depending on the stacking configuration [34, 45].
+Specifically, DFT calculations by Lee et al.
+[34] indicate that the AL-stacking configuration is energetically the most favorable, with the L-stacking only slightly higher in energy.
+This small energy difference suggests that a stacking rearrangement from AL to L can be induced by a moderate external perturbation such as a current pulse.
+To uncover the nature of this transformation, we examined the out-of-plane band dispersion at binding energies where differences between metallic and insulating regions are minimized, isolating the influence of interlayer effects.
+In the insulating regions, the kz dispersion shows a doubled periodicity and reduced bandwidth, consistent with interlayer dimerization and a doubling of the unit cell along the -axis, as previously reported [35].
+These features align with the AL-stacking configuration, in which vertical dimers of star-of-David clusters are stacked such that the central atom of one dimer aligns with one of the outer atoms of a neighboring cluster (See Fig.
+1B).
+The dashed white line in Fig.5B shows the calculated out-of-plane dispersion for this configuration, which matches well with our data from the insulating regions.
+In contrast, the metallic regions exhibit a single-period kz dispersion closely matching the primitive lattice periodicity and a broader bandwidth, indicating the absence of dimerization, in agreement with the DFT predictions for the L-configuration (See Fig.
+1B).
+These observations suggest that the current pulse disrupts the AL-stacked (dimerized) configuration, breaking the interlayer correlations that stabilize the insulating state.
+The system then transitions into the L-stacking configuration, a nearly degenerate configuration in energy, which restores a simpler stacking periodicity and enables coherent interlayer hopping, thereby giving rise to a three-dimensional metallic state.
+This hidden metallic phase represents an unusual electronic structure among transition metal dichalcogenides: the out-of-plane (interlayer) bandwidth exceeds the in-plane bandwidth.
+This is in agreement with transport measurements showing that the resistivity anisotropy in the NC-CDW state is close to one [46].
+Such an inversion of dimensionality, where transport and electronic coherence are dominated by the out-of-plane direction, challenges conventional views of quasi-2D physics and may open new avenues for engineering anisotropic transport and manipulating quantum phases in layered materials.
+Materials and Methods
+Sample preparation
+1T-TaS2 single crystal were grown as described in Ref.
+[47].
+Samples of 1T-TaS2 were prepared for nanoARPES experiments: Single crystals were cut into narrow bridges 200-500 in width.
 The samples were glued to a sapphire substrate and 4 contacts were made using Ag epoxy.
-The samples were cooled to 40K and cleaved under UHV conditions.
-The 4 contacts were used to drive a DC current pulse and to monitor the sample resistance.
-We emphasis that during the ARPES measurement no current was applied to the sample.
-IV.2 ARPES
+The samples were cooled to and cleaved under UHV conditions.
+The 4 contacts were used to drive a current pulse and to monitor the sample resistance in-situ.
+We emphasize that during the ARPES measurement no current was applied to the sample.
+ARPES measurements
 Spatially resolved ARPES experiments were carried out in the I05 beamline at Diamond light source, UK.
-The spatial resolution is set by the beam spot-size, which in our experiment was either 500nm using a Zone-plate or 5 micrometers using a capillary mirror.
+The spatial resolution is set by the beam spot-size, which in our experiment was either 500 nm using a Zone plate or 5 micrometers using a capillary mirror.
 Energy resolution was set to about 50 m eV.
@@
 The Fermi-level was determined by measuring a metallic sample at the same conditions.
-IV.3 calculation of k⟂
-We have used the three-step-model for the out-of-plane momentum k Damascelli (2004), where is the inner potential.
-We have used , obtained by following the normal-crystal periodicity over 3 BZs, demanding ().
-V Acknowledgments
+calculation of k⟂
+We have used the three-step-model for the out-of-plane momentum k [48], where is the inner potential. We have used , obtained by following the normal-crystal periodicity over 3 BZs, demanding ().
+Band fitting
+Peak fitting was done using Gaussian functions. When close to , the Gauassians were multiplied by the Fermi function, where was fixed to zero, and the temperature was set to .
+Acknowledgments
 We thank Assa Auerbach and Daniel Podolsky for useful discussions.
@@
 We acknowledge SOLEIL for provision of synchrotron radiation facilities under proposal 20230771 and we would like to thank Chiara Bigi and Francois Bertran for assistance in using beamline CASSIOPEE.
+Funding:
 The work at the Technion was supported by Israeli Science Foundation grant number ISF-1263/21.
+Author contributions:
+A.K.
+conceived the experiment.
+I.F.
+prepared the 1T-TaS2 single crystals.
+Y.N.
+performed the ARPES experiments and analyzed the ARPES data.
+R.A.G.K, I.M., A.D., N.R.
+S.K.P.
+and Y.J.
+performed ARPES measurements.
+A.L., M.D.W.
+and C.C.
+provided support at the synchrotron beam-line.
+Y.N.
+and A.K.
+wrote the manuscript with contributions from all authors.
+Competing interests:
+The authors declare that they have no competing interests.
+Data availability:
+All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Materials.
+Supplementary materials
+Figures S1 to S3